# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v4

@@ 5.3.1. Dataset Construction
Since SQL-Factory is primarily designed to synthesize executable SQL queries for given database schema from scratch, direct application to Text-to-SQL tasks requires transforming these SQLs into natural language (NL) descriptions. This enables the construction of ¡NL, SQL¿ training pairs for supervised learning.
Instead of generating SQL from NL, which often suffers from semantic mismatch and low exact match accuracy, we adopt a reverse formulation: converting SQL to NL. Recent studies (Wu et al., 2021; Guo et al., 2018; Zhang et al., 2023) have shown that LLMs are more reliable in generating faithful natural language descriptions for given SQL. This makes the SQL-to-NL approach a robust and efficient means for constructing high-quality data. To this end, we develop a multi-stage pipeline for SQL-to-NL conversion with consistency checking and optional SQL refinement. The steps are as follows:
(1) Similar Example Retrieval. To improve translation quality and ensure semantic alignment, we first retrieve top-$k$ similar examples (with $k$=5 in the experiment) from training set of BIRD or Spider using Retrieval Tool proposed in Section 4.1.2. These examples are used to construct the in-context prompt to teach model how to write fluent, semantically faithful questions from SQL. This retrieval-augmented generation ensures structural and linguistic alignment between training and inference distributions.
@@
(3) Semantic Consistency Verification. To further ensure the correctness of the constructed ¡NL, SQL¿ pairs, we introduce an LLM-based verifier to perform semantic consistency checking. This component is critical for filtering out misaligned instances where the generated natural language question does not faithfully reflect the intent or logic of the corresponding SQL query. Specifically, the verifier takes as input the SQL query and its corresponding natural language question, and is prompted to return a binary decision indicating whether the pair is semantically aligned. To improve reasoning quality and reduce false positives, we adopt a Chain-of-Thought (CoT) prompting strategy. Only those ¡NL, SQL¿ pairs that pass this verification step are retained, ensuring high semantic fidelity in the final dataset.
(4) SQL Simplification. In case where the SQL query contains redundant columns or sub-clauses not reflected in the NL intent, we use a follow-up prompt to refine the SQL accordingly. This ensures alignment between the question intent and query logic. All simplified SQLs are passed again through the verifier to ensure final consistency.
@@ 5.3.2. Model Training
We fine-tune two backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.3.3. Result and Analysis
To ensure fairness, we adopt the official prompt templates for each model. For codes models, we use the released prompt from the original repository but remove schema linking and value retrieval modules to match our setup. For all Qwen-family models, we adopt the default instruction prompt used by Qwen2.5-Coder (Hui et al., 2024). Each model is evaluated on two benchmarks: Spider and BIRD and the evaluation metrics include both executability and execution accuracy.
The results are shown in Table 3. After fine-tuning with the SQL-Factory dataset, both Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B exhibit substantial performance improvements. On Spider, Qwen2.5-Coder-1.5B+SQL-Factory achieves an executability of 97.20% and accuracy of 76.50%, while Qwen2.5-Coder-3B+SQL-Factory reaches 97.29% executability and 79.69% accuracy. Both outperform their corresponding base models by a large margin, demonstrate that SQL-Factory effectively boosts model performance even without increasing model size. On BIRD, which contains more complex and compositional queries, fine-tuned models again show superior performance. Qwen2.5-Coder-3B+SQL-Factory achieves 88.07% in executability and 46.09% in accuracy, clearly outperforming all baseline models of similar scale. These improvements are especially pronounced in the “Moderate” and “Challenging” categories, confirming the effectiveness of SQL-Factory in providing structurally rich and semantically aligned training data.
Overall, these findings show that with access to SQL-Factory, small and mid-sized models can achieve strong performance in Text-to-SQL tasks. This highlights SQL-Factory’s ability to deliver diverse and executable training examples that bridge the performance gap without requiring massive model sizes.
@@ 5.4. Query Clustering Task
Methods IIT Bombay UB Exam PocketData
Aouiche (Aouiche et al., 2006) 0.577 0.923 0.893
Aligon (Aligon et al., 2014) 0.535 0.799 0.898
Makiyama (Makiyama et al., 2015) 0.665 0.897 0.879
BERT-Base (Devlin et al., 2019) 0.524 0.769 0.773
BERT-Large (Devlin et al., 2019) 0.637 0.846 0.777
PreQR (Tang et al., 2022) 0.387 0.622 0.710
BERT-Base+SQL-Factory 0.376 0.558 0.656
BERT-Large+SQL-Factory 0.350 0.549 0.638
@@
We fine-tune both BERT-Base and BERT-Large models using all SQL queries synthesized by SQL-Factory, aiming to enhance their capability in capturing SQL-level semantics and improve performance in downstream clustering tasks. The training process employs contrastive learning, a method that leverages positive and negative sample pairs to enhance model understanding. Positive pairs are selected via the Retrieval Tool introduced in Section 4.1.2. Specifically, for each anchor query, we retrieve its top-20 most similar queries based on our hybrid similarity metric and treat them as positive samples. All remaining queries within the same batch serve as negative samples. This in-batch negative sampling strategy ensures scalability and efficiency, as it avoids the need to explicitly mine hard negatives from the full dataset while still providing sufficient contrastive signal.
@@
Table 4 presents the BetaCV scores across all models and benchmarks. Traditional rule-based methods such as Aouiche, Aligon, and Makiyama rely on handcrafted SQL features or syntax trees, and often fail to capture deep semantic similarities across structurally different but semantically related queries. PreQR improves upon this by leveraging a pretrained language model, but it lacks task-specific supervision and may underfit domain-specific patterns. Meanwhile, untuned BERT models benefit from general semantic priors but are still agnostic to SQL-specific structure and vocabulary.
By contrast, our contrastively fine-tuned BERT models significantly outperform all baselines. Both versions of the BERT model have achieved the best performance on these benchmarks after training. These results validate the effectiveness of SQL-Factory in producing high-quality, semantically diverse SQL queries that can be used as training signals for representation learning.

### v5

@@ 5.3.1. Dataset Construction
Since SQL-Factory is designed to synthesize executable queries for a given database schema from scratch, applying it to Text-to-SQL tasks requires transforming these SQLs into natural language (NL) descriptions. This enables the construction of ¡NL, SQL¿ training pairs for supervised learning.
Instead of generating SQL from NL, which often suffers from semantic mismatch and low exact match accuracy, we adopt a reverse formulation: converting SQL to NL. Recent studies (Wu et al., 2021; Guo et al., 2018; Zhang et al., 2023; Kobayashi et al., 2025) have shown that LLMs are more reliable in generating faithful natural language descriptions for given SQL, making SQL-to-NL a robust and efficient method for constructing high-quality datasets. To this end, we develop a multi-stage pipeline for SQL-to-NL conversion with consistency checking and optional SQL refinement. The steps are as follows:
(1) Similar Example Retrieval. To improve translation quality and ensure semantic alignment, we first retrieve top-$k$ similar examples (with kk=5 in the experiment) from training set of BIRD or Spider using Retrieval Tool proposed in Section 4.1.2. These examples are used to construct the in-context prompt to teach model how to write fluent, semantically faithful questions from SQL. This retrieval-augmented generation ensures structural and linguistic alignment between training and inference distributions.
@@
(3) Semantic Consistency Verification. To ensure the correctness of the constructed ¡NL, SQL¿ pairs, we introduce an LLM-based verifier to perform semantic consistency checking. This component filters out instances where the generated natural language question does not accurately reflect the intent or logic of the SQL query. While LLMs are not perfect in complex generative tasks like Text-to-SQL, recent work (Pourreza and Rafiei, 2023; Wang et al., 2025b; Chen et al., 2023a) shows they are highly reliable for binary classification and entailment-style tasks, supporting our choice of an LLM-based verifier. Specifically, the verifier takes as input the ¡NL, SQL¿ pair, and is prompted to return a binary decision indicating whether the pair is semantically aligned. To further improve reasoning quality, we adopt a Chain-of-Thought (CoT) prompting strategy, where the model generates reasoning steps in a step-by-step manner before producing the final decision. Only pairs passing this CoT-guided verification step are retained. In addition, we conducted a human-in-the-loop validation by randomly sampling 1,000 pairs and recruited five PhD-level STEM students with SQL experience to perform cross-annotation. Results show that only 4.6% of the samples were judged to be semantically inconsistent, indicating that the final dataset maintains high semantic fidelity and practical reliability.
@@ 5.3.2. Model Training
We fine-tune all the backbone models mentioned above, including Qwen2.5-Coder-1.5B/3B, LLaMA3.2-1B/3B, and Deepseek-Coder-1.3B, on the constructed dataset of ¡NL, SQL¿ pairs. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.3.3. Result and Analysis
To ensure fairness, we adopt the official prompt templates for each model. Each model is evaluated on the Spider and BIRD benchmarks, using both executability and execution accuracy as metrics. The results in Table 4 show that across all model scales and architectures, fine-tuning with SQL-Factory consistently improves performance on both benchmarks.
On Spider, Qwen2.5-Coder-1.5B improves in accuracy from 69.2% to 76.5%, while on BIRD, Deepseek-Coder-1.3B and LLaMA3.2-1B achieve relative accuracy gains exceeding 15–20 points. Similar trends are observed for executability. These gains are larger than those obtained from fine-tuning on the official datasets, indicating that SQL-Factory produces high-quality, executable training data that rivals or surpasses expert-curated sets. These results demonstrate that data generated by SQL-Factory significantly boosts model performance. The improvements are particularly evident in structurally complex cases, indicating the effectiveness of our schema-aware generation and expansion process.
@@ 5.4. Query Clustering Task
Method IIT Bombay UB Exam PocketData
Aouiche 0.577 0.923 0.893
Aligon 0.535 0.799 0.898
Makiyama 0.665 0.897 0.879
PreQR 0.387 0.622 0.710
BERT-Base 0.524 0.769 0.773
  +SQL-Factory 0.376 ($\downarrow$0.148) 0.558 ($\downarrow$0.211) 0.656 ($\downarrow$0.117)
BERT-Large 0.637 0.846 0.777
  +SQL-Factory 0.350 ($\downarrow$0.287) 0.549 ($\downarrow$0.297) 0.638 ($\downarrow$0.139)
@@
We fine-tune both BERT-Base and BERT-Large models using all SQL queries synthesized by SQL-Factory, aiming to enhance their capability in capturing SQL-level semantics. The training process employs contrastive learning, a method that leverages positive and negative sample pairs to enhance model understanding. Positive pairs are selected via the Retrieval Tool introduced in Section 4.1.2. Specifically, for each anchor query, we retrieve its top-20 most similar queries based on our hybrid similarity metric and treat them as positive samples. All remaining queries within the same batch serve as negative samples. This in-batch negative sampling strategy ensures scalability and efficiency, as it avoids the need to explicitly mine hard negatives from the full dataset while still providing sufficient contrastive signal.
@@
Table 5 presents the BetaCV scores across all models and benchmarks. Traditional rule-based methods such as Aouiche, Aligon, and Makiyama rely on handcrafted SQL features or syntax trees, and often fail to capture deep semantic similarities across structurally different but semantically related queries. PreQR improves upon this by leveraging a pretrained language model, but it lacks task-specific supervision and may underfit domain-specific patterns. Meanwhile, untuned BERT models benefit from general semantic priors but are still agnostic to SQL-specific structure and vocabulary. By contrast, our contrastively fine-tuned BERT models significantly outperform all baselines. Both versions of the BERT model have achieved the best performance on these benchmarks after training. These results validate the effectiveness of SQL-Factory in producing high-quality, semantically diverse SQL queries that can be used as training signals for representation learning.